How long would it take an investment of $12,225 at 3 3/4% interest compounded quarterly to triple?
Answer by stanbon(75887)

How long would it take an investment of $12,225 at 3 3/4% interest compounded quarterly to triple?
-----
A(t)=Ao(1+(r/n))^(nt)
------
3(12225)= 12225(1+(0.0375/4))^(4t)
(1+0.0094)^(4t) = 3
Take the log of both sides and solve for "t":
(4t) = log(3)/log(1.0094)
---
4t = 117.4221
---
t = 29.35
Rounding up you get: t = 30 years
=====================================
Cheers,
Stan H.
===========